# Positive formula for the product of conjugacy classes on the unitary group

### v1

@@ 4.2. Edge, vertex and face partitions
Let $i$ be the type of $B$, and suppose by contradiction that there is an edge $e^{0}$ of type $i$ or $i-1$ inside the strip $S_{B}$ which is contained in a strip $S_{\tilde{B}}$ for some $\tilde{B}\in\mathcal{P}_{e}$. Suppose first that $\tilde{B}\not=\{e^{0}\}$. Then, since $e^{0}$ shares a vertex with at least three edges of type $i+1$ colored 00, $S_{\tilde{B}}$ must contain of those three edges, called $f^{1}$; since $\mathcal{P}_{e}$ is non-crossing and $f^{1}\in\mathcal{E}\cap S_{\tilde{B}}$, $f^{1}\in\tilde{B}$ and thus $\tilde{B}$ is of type $i+1$. Since $e^{0}\in S_{\tilde{B}}$ and $\tilde{B}$ is of type $i+1$, the other edge $f^{2}$ of $S_{B}$ with same height as the one $f^{1}$ must also belong to $S_{\tilde{B}}$, and by the non-crossing condition we have $\tilde{B}=\{f^{1},f^{2}\}$. This contradicts the fact that two elements of the same block of $\mathcal{P}_{v}$ do not belong to the same block of $\mathcal{P}_{e}$.
If $e^{0}$ is a singleton of $\mathcal{P}_{e}$, then $e\in\mathcal{E}$ and belongs to at least two blocks of $\mathcal{P}_{v}$. Hence, there must be another edge $e^{1}$ between $e^{0}$ and $e$ in the strip $S_{B}$ which belongs to $\mathcal{E}$. Iterating the process yields an edge $\tilde{e}$ such that $e$ and $\tilde{e}$ belong to a same block $\tilde{B}$ in $\mathcal{P}_{v}$. Then, $\tilde{e}$ cannot be of type $i-1$, for otherwise this block would be a block of order $3$, contradicting the fact that the edge $f$ of type $i+1$ with $f_{i}=e_{i},f_{i+1}=e_{i+1}-1$, which would then belong to this block $\tilde{B}$, does not belong to $\mathcal{E}$. Similarly, if $\tilde{e}$ is of type $i$, then $e$ and $\tilde{e}$ would be boundary edges of type I piece with $2(r^{\prime}+1)$ edges, with $1\leq r^{\prime}<r$. This is impossible since the boundary $\partial_{2}S_{\{e,\tilde{e}\}}$ contains at most one edge which is not in $\mathcal{E}$. Hence, no edge of type $i$ or $i-1$ inside $S_{B}$ belongs to some strip $S_{B}$ for $B\in\mathcal{P}_{e}$ and thus none of those edges has been colored $2$ in the previous labelling. Therefore, one can color all type $i$ edge in $S_{\tilde{B}}$ different from $e,e^{\prime}$ with the label $7$ and all type $i-1$ edge in $S_{\tilde{B}}$ with label $6$.
@@
Finally, color all remaining edge with the label $2$. One then check that the labels on the boundary of any triangle of the puzzle satisfy the conditions of Figure 8, so that the labelling of edges of $E_{n}$ yields a genuine puzzle $P$. It is then straightforward to check that $P_{v}=\mathcal{P}_{v}$ and $P_{e}=\mathcal{P}_{e}$. The map $\Phi$ is thus surjective.
@@ 5. Discrete two-colored dual hive model
In this section, we associate to each puzzle of size $N$ a two-colored hive in the same spirit as in [16]. Beware that because of the rigid crossing from Figure 9, the discrete hives won’t be actual hives as in [16] but rather a dual hive. Let us fix in this section the number $n$ of edges colored 00 or $1$ on one edge of the puzzles. This number is the same on each edge of the puzzle and is part of the boundary data of the puzzle. The two colored dual hive associater to a puzzle will then be a decoration of the triangular grid of $T_{n}$ instead of $T_{N}$.
Definition 5.1 (Discrete boundary).
The boundary $\partial T_{n}$ of the discrete triangular grid $T_{n}$ is the intersection of $T_{n}$ with the boundary of the triangle.
We denote by by $F_{n}$ the set of faces of $T_{n}$. Remark that the faces are all triangular, and we decompose $F_{n}$ as $F_{n}=F_{n}^{+}\cup F_{n}^{-}$ depending on whether the triangles are direct or reversed. For $e\in E_{n},f\in T_{n}$, we write $e\in f$ when $e$ is an edge on the boundary of $f$. We denote by $E_{n}^{(\ell)}$ the set of edges of type $\ell$ and all the notation introduced for $T_{N}$ are still valid for $T_{n}$. Any union of two triangles sharing an edge $e$ is called a lozenge, and $e$ is then called the middle edge of the lozenge.
Remark that the boundary $\partial T_{n}$ can be decomposed into three subgraphs $\partial T_{n}^{(i)}$, $0\leq i\leq 2$, where each subgraph $\partial T_{n}^{(i)}$ consists of edges of type $i$.
Definition 5.2 (Discrete two-colored dual hive).
A two-colored discrete dual hive of size $n$ is given by the following combinatorial data on $T_{n}$:
  • a color map $C:E_{n}\rightarrow\{0,1,3,m\}$, such that the boundary colors around each triangular face in the clockwise order is either $(0,0,0),(1,1,1),(1,0,3)$ or $(0,1,m)$ up to a cyclic rotation.
  • a label map $L:E_{n}\rightarrow\mathbb{N}$, with the two following conditions:
@@
      for all $f\in F_{n}$ with boundary edges $e^{0},e^{1},e^{2}$, $L(e^{0})+L(e^{1})+L(e^{2})=N-1$ except when $f\in F_{n}^{-}$ with boundary colors different from $\{0,1,m\}$, in which case $L(e^{0})+L(e^{1})+L(e_{2})=N-2$,
@@
        $L(e)\geq L(e^{\prime})$ if $e^{\prime}_{\ell+1}=e_{\ell+1}+1$ and none of the edge different of $e^{\prime}$ are colored $m$,
@@
        $L(e)>L(e^{\prime})$ if $e_{\ell}>e^{\prime}_{\ell}$ or if $e^{\prime}_{\ell+1}=e_{\ell+1}+1$ and one of the boundary edges different from $e^{\prime}$ is colored $m$.
The boundary value $[(c^{(0)},c^{(1)},c^{(2)}),(l^{(0)},l^{(1)},l^{(2)})]$ of a two-colored discrete dual hive is the restriction of $(C,L)$ to $\partial T_{n}$, where $c^{(i)}\in\{b,r,m^{-},m^{+}\}^{n}$ (resp. $l^{(i)}\in\mathbb{N}^{n}$) is the restriction of $C$ (resp. $L$) to $\partial T_{n}^{(i)}$, for $0\leq i\leq 2$.
@@
As a corollary of the Condition $(2)$ on the label map, we have $L(e)>L(e^{\prime})$ for any pair of edges $e,e^{\prime}$ of same type $\ell$ such that $e_{\ell}>e^{\prime}_{\ell}$ and $e_{\ell+1}\leq e^{\prime}_{\ell+1}$. Indeed, it suffices to show this for $e,e^{\prime}$ such that $e_{\ell}=e^{\prime}_{\ell}+1$ and $e^{\prime}_{\ell+1}\in\{e_{\ell+1},e_{\ell+1}+1\}$. The case $e^{\prime}_{\ell+1}=e_{\ell+1}$ is given by Condition $(2.c)$, and we now suppose that $e^{\prime}_{\ell+1}=e_{\ell+1}+1$. Let $e^{\prime\prime}$ be such that $e^{\prime\prime}_{\ell+1}=e^{\prime}_{\ell+1}$ and $e^{\prime\prime}_{\ell}=e_{\ell}+1$. If the middle edge of the lozenge with boundary $e^{\prime},e^{\prime\prime}$ is not colored $m$, then by Condition $(2.c)$ we have $L(e^{\prime\prime})>L(e^{\prime})$, and then by $(3.a)$ or $(3.b)$, we get $L(e)\geq L(e^{\prime\prime})>L(e^{\prime})$. If the middle edge of the lozenge with boundary $e^{\prime},e^{\prime\prime}$ is colored $m$, then $L(e^{\prime})=L(e^{\prime\prime})$. Then, the middle edge colored $m$ of this lozenge is then a boundary edge of the lozenge with boundary $e^{\prime\prime},e$ different from $e^{\prime\prime}$ and $e$, so that $(3.c)$ implies that $L(e)>L(e^{\prime\prime})=L(e^{\prime})$.
For $(c,l)=(c^{(0)},c^{(1)},c^{(2)},l^{(0)},l^{(1)},l^{(2)})\in\{b,r\}^{3n}\times\mathbb{N}^{3n}$, we denote by $H(c,l)$ the set of two-colored discrete dual hives with boundary value $(c,l)$. For any boundary date $(c,l)$, denote by $\omega^{c,l}$ the triple of words of length $N$ such that $\omega^{c,l}_{\ell}(l^{(\ell)}(i))=c^{(\ell)}(i)$ for $\ell\in\{0,1,2\}$ and $1\leq i\leq n$ and $\omega^{c,l}_{\ell}(j)=2$ for other $10\leq j\leq N-$ 1.
Proposition 5.4.
For any valid pair $(c,l)\in\{b,r\}^{3n}\times\mathbb{N}^{3n}$,
$|H(c,l)|=|P(\omega^{c,l})|.$
Remark that Proposition 5.8 actually gives more generally the expression of the structure coefficients of the two-step flag variety in terms of counting of integers points into a convex polytope.
We prove this proposition by exhibiting a bijection $\Phi:P(\omega^{c,l})\rightarrow H(c,l)$ which is of interest for the sequel. Given a puzzle $P\in P(\omega^{c,l})$, let $\mathcal{G}_{P}$ be the corresponding graph introduced in Section 4.3.
@@
Let us first transform the $\mathcal{G}_{P}$ into a new graph $\widehat{G}_{P}$ by blowing up each vertex $v\in\mathcal{P}_{v}$ of size $2(r+1)$ as follows. First, by Lemma 4.7, a vertex $v\in\mathcal{P}_{v}$ of size $2(r+1)$ has adjacent edges $(B^{1},\ldots,B^{2r+2})$ (indexed in the cyclic order around the vertex with firest edge colored $1$) such that $B^{1},B^{r+2}$ have type $\ell\in\{0,1,2\}$ and are colored $1$ and $B^{i},i\not\in\{1,r+2\}$ have type $\ell+1$ and are colored 00. Introduce $12r-$ 1 new edges $\tilde{B}^{1},\ldots,\tilde{B}^{2r-1}$ of type $\ell-1,\ell,\ldots,\ell-1$ and colored $m,1,\ldots,m$ and transform $v$ into $2r$ vertices $v^{1},\ldots,v^{2r}$ such that the edges adjacent to $v^{2j+1}$ are $(\tilde{B}^{2j},B^{2r+2-j},\tilde{B}^{2j+1})$ and edges adjacent to $v^{2j+2}$ are $(\tilde{B}^{2j+1},\tilde{B}^{2j+2},B^{j+1})$ (with the convention $\tilde{B}^{0}=B^{1}$ and $\tilde{B}^{2r}=B^{r+2}$. We define the height of $\tilde{B}^{i}$ as $h(\tilde{B}^{2i})=h(B^{1})$ and $h(\tilde{B}^{2i-1})=N-1-h(B^{1})-h(B^{2i})$ for $11\leq i\leq r-$ 1.
@@
The resulting graph has then only vertices of degree $3$ or singletons (which correspond to edges of $T_{N}$ on the boundary of the triangle).
Lemma 5.5.
Let $B_{f}$ be a face in $\widehat{G}_{P}$. Then, the the boundary of $B_{f}$ has
@@
The resulting planar graph $\widehat{\mathcal{G}}_{P}$ is thus a graph with only trivalent vertices and hexagonal inner faces. From each side of the triangle $T_{N}$, there are $n-2$ faces $B\in\mathcal{P}_{f}$ which have degree $4$ (one for each pair of consecutive boundary edge labeled 00 or $1$ on a same side of $T_{N}$) and from each extreme vertex of $T_{N}$ there is a face of degree $2$.
Let us denote by $\widetilde{\mathcal{G}}_{P}$ the dual graph, namely the graph whose vertices are faces of $\widehat{\mathcal{G}}_{P}$, faces are vertices of $\widehat{\mathcal{G}}_{P}$ and such that there is one edge between each neighboring faces of $\widetilde{\mathcal{G}}_{P}$ (which correspond then vertices of $\widehat{\mathcal{G}}_{P}$).
Lemma 5.6.
There is an isomorphism from $\widetilde{\mathcal{G}}_{P}$ to $T_{n}$ mapping edges of type $\ell$ of $\widetilde{\mathcal{G}}_{P}$ to edges of type $\ell$ of $T_{n}$.
@@
Since vertices of $\widehat{\mathcal{G}}_{P}$ are trivalent, faces of $\widetilde{\mathcal{G}}_{P}$ are triangular. Similarly, inner faces of $\widehat{\mathcal{G}}_{P}$ have degree $6$, and thus inner vertices of $\widetilde{\mathcal{G}}_{P}$ have degree $6$. Hence, $\widetilde{\mathcal{G}}_{P}$ is isomorphic to a polygon $H$ of the planar triangular grid.
Since the sequence of degrees of the $3n$ outer faces of $\widehat{\mathcal{G}}_{P}$ is
$(2,\underset{n\text{ times}}{\underbrace{4,\ldots,4}},2,\underset{n\text{ times}}{\underbrace{4,\ldots,4}},2,\underset{n\text{ times}}{\underbrace{4,\ldots,4}}),$
the same holds for the sequence of degrees of outer vertices of $\widetilde{\mathcal{G}}_{P}$. Remark that for each vertex of degree $4$ (resp. $2$) in $\widetilde{\mathcal{G}}_{P}$, the angle of the boundary at the corresponding vertex in $H$ is $\pi$ (resp. $35\pi/$ 3). We deduce that the boundary of the $\widetilde{\mathcal{G}}_{P}$ is isomorphic to the one of $T_{n}$, and thus $\widetilde{\mathcal{G}}_{P}$ is isomorphic to $T_{n}$. Let us denote by $\Phi$ the corresponding bijection between set of edges.
Remark that around every triangle of $\widetilde{\mathcal{G}}_{P}$ the type of the edges is $(\ell,\ell+1,\ell+2)$ (this is true for faces coming from trivalent vertex of $\mathcal{G}_{P}$ and true by construction for faces coming from the blowing up of higher order vertices of $\widehat{\mathcal{G}}_{P}$). We deduce that all edges with the same type in $\widetilde{\mathcal{G}}_{P}$ are sent through $\Phi$ to edges with the same orientation in $T_{n}$. Up to composing $\Phi$ with an internal rotational symmetry of $T_{n}$, we can thus assume that $\Phi$ preserves the type of the edges. ∎
Each edge $B$ of $\widetilde{\mathcal{G}}_{P}$ has then a color $c(B)$ and a height $h(B)$ coming from the dual edge of $\widehat{\mathcal{G}}_{P}$. Composing with $\Phi$, this yields maps $C:E_{n}\rightarrow\mathbb{N}$ and $L:E_{n}\rightarrow\mathbb{N}$ with $C=\Phi\circ c$ and $L=\Phi\circ h$.
Lemma 5.7.
The resulting pair of maps $(C,L)$ is a discrete two-colored dual hive.
@@
Lemma 4.7 and the he color rules introduced before in case of blowing-up of even degree vertices yield that edges around each trivalent vertex of $\widetilde{\mathcal{G}}_{P}$ are colored $(0,0,0)$, $(1,1,1)$, $(1,0,3)$ or $(0,1,m)$ in the clockwise order, which translate into the same color rule around each triangular face of $T_{n}$.
It remains to prove that the map $L$ on $E_{n}$ satisfies the two conditions of Definition 5.2. The sum condition (1) around a triangle is a consequence of Lemma 4.5 in case no edge is colored $m$, and the direct deduction of the blowing up of vertices of even degree in case one of the edges is colored $m$.

### v2

@@ 4.3. Edge, vertex and face partitions
Let $i$ be the type of $B$, and suppose by contradiction that there is an edge $e^{0}$ of type $i$ or $i-1$ inside the strip $S_{B}$ which is contained in a strip $S_{\tilde{B}}$ for some $\tilde{B}\in\mathcal{P}_{e}$. Suppose first that $\tilde{B}\not=\{e^{0}\}$. Then, since $e^{0}$ shares a vertex with at least three edges of type $i+1$ colored 00, $S_{\tilde{B}}$ must contain one of those three edges, called $f^{1}$; since $\mathcal{P}_{e}$ is non-crossing and $f^{1}\in\mathcal{E}\cap S_{\tilde{B}}$, $f^{1}\in\tilde{B}$ and thus $\tilde{B}$ is of type $i+1$. Since $e^{0}\in S_{\tilde{B}}$ and $\tilde{B}$ is of type $i+1$, the other edge $f^{2}$ of $S_{B}$ with same height as $f^{1}$ must also belong to $S_{\tilde{B}}$, and by the non-crossing condition we have $\tilde{B}=\{f^{1},f^{2}\}$. This contradicts the fact that two elements of the same block of $\mathcal{P}_{v}$ do not belong to the same block of $\mathcal{P}_{e}$.
If $e^{0}$ is a singleton of $\mathcal{P}_{e}$, then $e\in\mathcal{E}$ and belongs to at least two blocks of $\mathcal{P}_{v}$. Hence, there must be another edge $e^{1}$ between $e^{0}$ and $e$ in the strip $S_{B}$ which belongs to $\mathcal{E}$. Iterating the process yields an edge $\tilde{e}$ such that $e$ and $\tilde{e}$ belong to a same block $\tilde{B}$ in $\mathcal{P}_{v}$. Then, $\tilde{e}$ cannot be of type $i-1$ otherwise this block would be a block of order $3$, contradicting the fact that the edge $f$ of type $i+1$ with $f_{i}=e_{i},f_{i+1}=e_{i+1}-1$, which would then belong to this block $\tilde{B}$, does not belong to $\mathcal{E}$. Similarly, if $\tilde{e}$ is of type $i$, then $e$ and $\tilde{e}$ would be boundary edges of type I piece with $2(r^{\prime}+1)$ edges, with $1\leq r^{\prime}<r$. This is impossible since the boundary $\partial_{2}S_{\{e,\tilde{e}\}}$ contains at most one edge which is not in $\mathcal{E}$. Hence, no edge of type $i$ or $i-1$ inside $S_{B}$ belongs to some strip $S_{B}$ for $B\in\mathcal{P}_{e}$ and thus none of those edges has been colored $2$ in the previous labelling. Therefore, one can color all type $i$ edge in $S_{\tilde{B}}$ different from $e,e^{\prime}$ with the label $7$ and all type $i-1$ edge in $S_{\tilde{B}}$ with label $6$.
@@
Color all remaining edges with the label $2$. One then checks that the labels on the boundary of any triangle of the puzzle satisfy the conditions of Figure 8, so that the labelling of edges of $E_{N}$ yields a genuine puzzle $P$. It is then straightforward to check that $P_{v}=\mathcal{P}_{v}$ and $P_{e}=\mathcal{P}_{e}$. The map $\Phi$ is thus surjective.
@@ 5. Discrete two-colored dual hive model
In this section, we associate to each puzzle of size $N$ a two-colored hive in the same spirit as in [19]. Beware that because of the rigid crossings from Figure 9, the discrete hives won’t be actual hives as in [19] but rather a dual hive. Let us fix in this section the number $k$ of edges colored 00 or $1$ on one edge of the puzzles. This number is the same on each edge of the puzzle and is part of the boundary data of the puzzle. The two colored dual hive associated to a puzzle will then be a decoration of the triangular grid $T_{k}$ instead of $T_{N}$. All the notation introduced for $T_{N}$ are thus still valid for $T_{k}$.
Definition 5.1 (Two-colored discrete dual hive).
A two-colored discrete dual hive of size $(k,N)$ is given by the following combinatorial data on $T_{k}$:
  • a color map $C:E_{k}\rightarrow\{0,1,3,m\}$, such that the boundary colors around each triangular face in the clockwise order is either $(0,0,0),(1,1,1),(1,0,3)$ or $(0,1,m)$ up to a cyclic rotation.
  • a label map $L:E_{k}\rightarrow\mathbb{N}$, with the two following conditions:
@@
      for all $f\in F_{k}$ with boundary edges $e^{0},e^{1},e^{2}$, $L(e^{0})+L(e^{1})+L(e^{2})=N-1$ except when $f\in F_{k}^{-}$ with boundary colors different from $\{0,1,m\}$, in which case $L(e^{0})+L(e^{1})+L(e^{2})=N-2$,
@@
        $L(e)\geq L(e^{\prime})$ if $e^{\prime}_{\ell+1}=e_{\ell+1}+1$ and no edge different from $e^{\prime}$ is colored $m$,
@@
        $L(e)>L(e^{\prime})$ if either $e_{\ell}>e^{\prime}_{\ell}$ or if both $e^{\prime}_{\ell+1}=e_{\ell+1}+1$ and one of the boundary edges different from $e^{\prime}$ is colored $m$.
The boundary value $[(c^{(0)},c^{(1)},c^{(2)}),(l^{(0)},l^{(1)},l^{(2)})]$ of a two-colored discrete dual hive is the restriction of $(C,L)$ to $\partial T_{k}$, where $c^{(i)}\in\{0,1,3,m\}^{k}$ (resp. $l^{(i)}\in\mathbb{N}^{k}$) is the restriction of $C$ (resp. $L$) to $\partial T_{k}^{(i)}$, for $0\leq i\leq 2$.
For $(c,l)=(c^{(0)},c^{(1)},c^{(2)},l^{(0)},l^{(1)},l^{(2)})\in\{0,1\}^{3k}\times\mathbb{N}^{3k}$, we denote by $\mathcal{H}(c,l,N)$ the set of two-colored discrete dual hives with boundary value $(c,l)$.
@@
As a corollary of the Condition $(2)$ on the label map, we have $L(e)>L(e^{\prime})$ for any pair of edges $e,e^{\prime}$ of same type $\ell$ such that $e_{\ell}>e^{\prime}_{\ell}$ and $e_{\ell+1}\leq e^{\prime}_{\ell+1}$. Indeed, it suffices to show this for $e,e^{\prime}$ such that $e_{\ell}=e^{\prime}_{\ell}+1$ and $e^{\prime}_{\ell+1}\in\{e_{\ell+1},e_{\ell+1}+1\}$. The case $e^{\prime}_{\ell+1}=e_{\ell+1}$ is given by Condition $(2.c)$, and we now suppose that $e^{\prime}_{\ell+1}=e_{\ell+1}+1$. Let $e^{\prime\prime}$ be such that $e^{\prime\prime}_{\ell+1}=e^{\prime}_{\ell+1}$ and $e^{\prime\prime}_{\ell}=e_{\ell}=e_{\ell}^{\prime}+1$. If the middle edge of the lozenge with boundary $e^{\prime},e^{\prime\prime}$ is not colored $m$, then by Condition $(2.c)$ we have $L(e^{\prime\prime})>L(e^{\prime})$, and then by $(2.a)$ or $(2.b)$, we get $L(e)\geq L(e^{\prime\prime})>L(e^{\prime})$. If the middle edge of the lozenge with boundary $e^{\prime},e^{\prime\prime}$ is colored $m$, then $L(e^{\prime})=L(e^{\prime\prime})$. Then, the middle edge colored $m$ of this lozenge is then a boundary edge of the lozenge with boundary $e^{\prime\prime},e$ different from $e^{\prime\prime}$ and $e$, so that $(2.c)$ implies that $L(e)>L(e^{\prime\prime})=L(e^{\prime})$.
For any triple $\omega=(\omega_{0},\omega_{1},\omega_{2})$ of words $\{0,1,2\}^{N}$ with $k_{0}$ occurrences of 00 and $k_{1}$ occurrences of $1$, denote by $(c(\omega),l(\omega))$ the sequence $(c^{(0)},c^{(1)},c^{(2)},l^{(0)},l^{(1)},l^{(2)})\in\{0,1\}^{3k}\times\mathbb{N}^{3k}$ where $k=k_{0}+k_{1}$ and $c^{(i)}$ is the word obtained from $\omega_{i}$ by deleting the letters $2$ and $l^{(i)}$ is the sequence of positions of the letters 00 or $1$ in $\omega_{i}$. The following result gives a formulation of Theorem 4.5 in terms of integer points counting of polytopes.
Theorem 5.3 (Dual hive in the two-step case).
For any triple $\omega=(\omega_{0},\omega_{1},\omega_{2})$ of words $\{0,1,2\}^{N}$ with $k_{0}$ occurrences of 00 and $k_{1}$ occurrences of $1$,
(5.1) $\langle\sigma_{\omega_{0}}\sigma_{\omega_{1}}\sigma_{\omega_{2}},\sigma_{0}\rangle_{H\mathbb{F}(k_{0},k_{1},N)}=\#\mathcal{H}(c(\omega),l(\omega),N).$
By Corollary 4.6, the latter theorem directly yields a similar expression of the quantum Littlewood-Richardson coefficients as the number of integer points in discrete dual hives. For three partitions $\lambda,\mu,\nu$ of length $n$ with first part smaller than $N-n$ and such that $|\lambda|+|\mu|=|\nu|+dN$, set
$H(\lambda,\mu,\nu,N)=\mathcal{H}(c(\omega),l(\omega),N),$
where $\omega$ is the triple of words in $\{0,1,2\}^{N}$ built as in Corollary 4.6 for $\lambda^{1}=\lambda,\,\lambda^{2}=\mu,\,\lambda^{0}=\nu$ and for the corresponding $d$.
Corollary 5.4 (Dual hive for the $q$-LR coefficients).
For $\lambda,\mu,\nu$ of length $n$ with first part smaller than $N-n$ and such that $|\lambda|+|\mu|=|\nu|+dN$,
$c_{\lambda,\mu}^{\nu,d}=\#H(\lambda,\mu,\nu,N).$
The rest of this section is then devoted to a proof of Theorem 5.3, which is obtained by exhibiting a bijection $\zeta:P(\omega)\rightarrow\mathcal{H}(c(\omega),l(\omega),N)$.
Given a puzzle $P\in P(\omega)$, let $\mathcal{G}_{P}$ be the corresponding graph introduced in Section 4.4. Let us first transform the graph $\mathcal{G}_{P}$ into a new graph $\widehat{G}_{P}$ by blowing up each vertex $v\in\mathcal{P}_{v}$ of size $2(r+1)$ as follows.
Definition 5.5 (Blowup of vertex).
@@
Let vertex $v\in\mathcal{P}_{v}$ be a vertex of size $2(r+1)$ with adjacent edges $(B^{1},\ldots,B^{2r+2})$ (indexed in the cyclic order) such that $B^{1},B^{r+2}$ have type $\ell\in\{0,1,2\}$ and are colored $1$ and $B^{i},i\not\in\{1,r+2\}$ have type $\ell+1$ and are colored 00. Introduce $12r-$ 1 new edges $\tilde{B}^{1},\ldots,\tilde{B}^{2r-1}$ of type $\ell-1,\ell,\ldots,\ell-1$ and colored $m,1,\ldots,m$ and transform $v$ into $2r$ vertices $v^{1},\ldots,v^{2r}$ such that the edges adjacent to $v^{2j+1}$ are $(\tilde{B}^{2j},B^{2r+2-j},\tilde{B}^{2j+1})$ and edges adjacent to $v^{2j+2}$ are $(\tilde{B}^{2j+1},\tilde{B}^{2j+2},B^{j+2})$ with the convention $\tilde{B}^{0}=B^{1}$ and $\tilde{B}^{2r}=B^{r+2}$. We define the height of $\tilde{B}^{i}$ as $h(\tilde{B}^{2i})=h(B^{1})$ and $h(\tilde{B}^{2i-1})=N-1-h(B^{1})-h(B^{2i})$ for $11\leq i\leq r-$ 1.
The resulting graph is called the blowup of $v$.
The picture of the blowup of a vertex of size $6$ is given in Figure 11.
$v$
$v_{1}$
$v_{2}$
$v_{3}$
$v_{4}$
$B_{1}$
$B_{1}$
$B_{2}$
$B_{2}$
$B_{3}$
$B_{3}$
$B_{4}$
$B_{4}$
$B_{5}$
$B_{5}$
$B_{6}$
$B_{6}$
$\tilde{B}_{1}$
$\tilde{B}_{2}$
$\tilde{B}_{3}$
Figure 11. Blowup of a vertex of size $6$ : edges in blue (resp. red, resp. green) are of type $\ell$ (resp. $\ell+1$, resp. $\ell+2$) and colored $1$ (resp. 00, resp. $m$).
Definition 5.6 (Blowup of the graph of a puzzle).
The blowup $\widehat{G}_{P}$ of the graph $\mathcal{G}_{P}$ is the graph obtained by blowing up every vertex of size $2(r+1)$, $r\geq 1$.
@@
Remark that the blowup of graph is well-defined because any vertex of size $2(r+1)$ of $\mathcal{G}_{P}$ has the form of Definition 5.5 thanks to Lemma 4.9. The blowup graph has then only vertices of degree $3$ or singletons (which correspond to edges of $T_{N}$ on the boundary of the triangle).
Lemma 5.7 (Faces in $\widehat{G}_{P}$).
Let $B_{f}$ be a face in $\widehat{G}_{P}$. Then, the boundary of $B_{f}$ has
@@
The resulting planar graph $\widehat{\mathcal{G}}_{P}$ is thus a graph with only trivalent vertices and hexagonal inner faces. From each side of the triangle $T_{N}$, there are $k-2$ faces $B\in\mathcal{P}_{f}$ which have degree $4$ (one for each pair of consecutive boundary edge labeled 00 or $1$ on a same side of $T_{N}$) and from each extreme vertex of $T_{N}$ there is a face of degree $2$.
Let us denote by $\widetilde{\mathcal{G}}_{P}$ the dual graph, namely the graph whose vertices are faces of $\widehat{\mathcal{G}}_{P}$, faces are vertices of $\widehat{\mathcal{G}}_{P}$ and such that there is one edge between each neighboring faces of $\widetilde{\mathcal{G}}_{P}$ (which correspond then to vertices of $\widehat{\mathcal{G}}_{P}$).
Lemma 5.8 (Dual graph to $T_{k}$).
There is an isomorphism from $\widetilde{\mathcal{G}}_{P}$ to $T_{k}$ mapping edges of type $\ell$ of $\widetilde{\mathcal{G}}_{P}$ to edges of type $\ell$ of $T_{k}$.
@@
Since vertices of $\widehat{\mathcal{G}}_{P}$ are trivalent, faces of $\widetilde{\mathcal{G}}_{P}$ are triangular. Similarly, inner faces of $\widehat{\mathcal{G}}_{P}$ have degree $6$, and thus inner vertices of $\widetilde{\mathcal{G}}_{P}$ have degree $6$. Hence, $\widetilde{\mathcal{G}}_{P}$ is isomorphic to a polygon $H$ of the planar triangular grid. Since the sequence of degrees of the $3k$ outer faces of $\widehat{\mathcal{G}}_{P}$ is
$(2,\underset{k\text{ times}}{\underbrace{4,\ldots,4}},2,\underset{k\text{ times}}{\underbrace{4,\ldots,4}},2,\underset{k\text{ times}}{\underbrace{4,\ldots,4}}),$
the same holds for the sequence of degrees of outer vertices of $\widetilde{\mathcal{G}}_{P}$. Remark that for each vertex of degree $4$ (resp. $2$) in $\widetilde{\mathcal{G}}_{P}$, the angle of the boundary at the corresponding vertex in $H$ is $\pi$ (resp. $35\pi/$ 3). We deduce that the boundary of the $\widetilde{\mathcal{G}}_{P}$ is isomorphic to the one of $T_{k}$, and thus $\widetilde{\mathcal{G}}_{P}$ is isomorphic to $T_{k}$. Let us denote by $\zeta:E(\widetilde{\mathcal{G}}_{P})\rightarrow E_{k}$ the corresponding bijection between set of edges.
Remark that around every triangle of $\widetilde{\mathcal{G}}_{P}$ the type of the edges is $(\ell,\ell+1,\ell+2)$ (this is true for faces coming from trivalent vertex of $\mathcal{G}_{P}$ and true by construction for faces coming from the blowing up of higher order vertices of $\widehat{\mathcal{G}}_{P}$). We deduce that all edges with the same type in $\widetilde{\mathcal{G}}_{P}$ are sent through $\zeta$ to edges with the same orientation in $T_{k}$. Up to composing $\zeta$ with an internal rotational symmetry of $T_{k}$, we can thus assume that $\zeta$ preserves the type of the edges. ∎
Each edge $B$ of $\widetilde{\mathcal{G}}_{P}$ has then a color $c(B)$ and a height $h(B)$ coming from the dual edge of $\widehat{\mathcal{G}}_{P}$. Composing with $\zeta^{-1}$ yields maps $C:E_{k}\rightarrow\mathbb{N}$ and $L:E_{k}\rightarrow\mathbb{N}$ with $C=c\circ\zeta^{-1}$ and $L=h\circ\zeta^{-1}$.
Lemma 5.9 (Image is a dual hive).
The resulting pair of maps $(C,L)$ is a discrete two-colored dual hive $\zeta(P)$ in $\mathcal{H}(c(\omega),l(\omega),N)$.
@@
Lemma 4.9 and the color rules introduced before in case of blowing-up of even degree vertices yield that edges around each trivalent vertex of $\widetilde{\mathcal{G}}_{P}$ are colored $(0,0,0)$, $(1,1,1)$, $(1,0,3)$ or $(0,1,m)$ in the clockwise order, which translates into the same color rule around each triangular face of $T_{k}$.
It remains to prove that the map $L$ on $E_{k}$ satisfies the two conditions of Definition 5.1. The sum condition (1) around a triangle is a consequence of Lemma 4.7 in case no edge is colored $m$, and the direct deduction of the blowing up of vertices of even degree in case one of the edges is colored $m$.